OpenStudy (anonymous):

i need to find dy/dt when x=8 and dx/dt=10 for xy=4

OpenStudy (anonymous):

\[xy=4\] and i guess \(x=x(t)\) and \(y=y(t)\) i.e. they are both functions of \(t\) so via the product rule you get \[x'(t)y+y'(t)x=0\]

OpenStudy (anonymous):

it is easy to write \(x'\) than \(\frac{dx}{dt}\)

OpenStudy (anonymous):

now if \(x=8\) then since \(xy=4\) you know \(y=\frac{1}{2}\)

OpenStudy (anonymous):

and you are told \(x'=10\)

OpenStudy (anonymous):

solve \[x'(t)y+y'(t)x=0\] which is now \[10\times \frac{1}{2}+y'\times 8=0\] for \(y'\)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw